For a continuous random variable X, P(20 ≤ X ≤ 66) = 0.31 and P(X > 66) = 0.25. Calculate the following probabilities. (Leave no cells blank - be certain to enter "0" wherever required. Round your answers to 2 decimal places.)
a. P(X < 66)
b. P(X < 20)
c. P(X = 66)

Answer by greenestamps(13198)



For a continuous random variable, the probability that the value is exactly any one number is zero.

c. P(X = 66) = 0

a. Given that P(X > 66) = 0.25, and now knowing that P(X = 66) is 0, P(X < 66) = 1-0.25 = 0.75.

b. Given that P(20 <= X <= 66) = 0.31, and now knowing that P(X < 66) = 0.75, P(X < 20) = 0.75-0.31 = 0.44.